# Accurately simulating core-collapse self-interacting dark matter halos

### v2

@@ 4.3 Angular and velocity dependence of the cross-section
Figure 15: Mass enclosed within 10 pc as a function of time for the satellite halos. The upper panel gives the results for a velocity-independent cross-section (simulations U, V, W, and X), and the lower panel is for velocity-dependent cross-sections (simulations Y and Z). More details on the simulations can be found in Tab. 1. In addition, the grey dotted line indicates the preferred mass range for the perturber of the stellar stream GD-1 (Bonaca et al., 2019).
In this last part, we investigate the role of the velocity and angular dependence of the self-interaction for the evolution of the satellite halo. As for the isolated halo, we employ isotropic and forward-dominated cross-sections following the velocity dependence given by Eq. (4) or being velocity-independent. In detail, we employ the same cross-section as in Sect. 3.8. The results for the enclosed mass within 10 pc are shown in Fig. 15.
@@
Interestingly, we find that the angular dependence starts to play a role in the velocity-independent case, although the scatterings between the DM of the host and the satellite are neglected. The small-angle scattering leads here to a longer collapse time compared to the isotropic cross-section (Fig. 15). For the isolated halo, the two cross-sections resulted in the same collapse time (Fig. 12). Fischer et al. (2021a) showed that the thermalisation process depends on the angular dependence of the cross-section, i.e. large-angle scattering populates the high-velocity tail of the Maxwell-Boltzmann distribution faster compared to small-angle scattering. In combination with tidal stripping, this allows explaining the different collapse times. Moreover, the difference arising from the angular dependence is minor in the velocity-dependent case where scattering with a high relative velocity is strongly suppressed (bottom panel of Fig. 15).
@@ 4.4 Fit with King model
Figure 16: Density profile fitted by a King model for a satellite halo. The results are for the simulation with a velocity-dependent isotropic cross-section. All simulation parameters are given in Tab. 1 under simulation Y.
@@
With this, we follow Zhang et al. (2025) who also used a King profile to describe the collapse phase of a potential GD-1 perturber, motivated by research on star clusters. To determine the density and radial parameters ($\rho_{0}$, $r_{\mathrm{c}}$), we use a limited radial range with $r<0.5\,\mathrm{kpc}$ only. For the fit, we maximise a likelihood based on Poisson statistics analogously to the description in Sect. 4 by Fischer et al. (2021a).
In Fig. 16, we show the fit of the King model for the collapse phase using our simulation with a velocity-dependent isotropic cross-section (simulation Y according to Tab. 1). It becomes visible that the King model provides a reasonable fit to the inner region of a satellite halo undergoing core collapse. This is in line with the findings by Zhang et al. (2025) and may work well enough to fit observational data as discussed by the same authors. In addition, we fit the King mode to the isolated halo and show the density profile as well as the time evolution of the fitted parameter in Appendix D.
@@ 5 Discussion
In this section, we discuss the limitations of our work and of the employed numerical scheme. We mention physical processes that we have neglected and that would be worth including in the modelling, and we highlight directions to improve the modelling of collapsing SIDM halos.
@@
In our simulations, we have neglected scatterings between satellite and host particles. However, they can play a crucial role in shaping the evolution of the satellite. How strongly the DM of the satellite is affected depends on how strong the cross-section is on the relevant velocity scale, which is typically larger than the velocities within the satellite. Hence for a cross-section that is decreasing as a function of velocity this is less important than for a velocity-independent cross-section (e.g. Silverman et al., 2022; Zeng et al., 2022, 2025). In addition, the angular dependence of the cross-section can play an important role. Satellite halos are strongly affected if the cross-section is forward-dominated (Fischer et al., 2022) and the effect becomes weaker for velocity-dependent cross-sections (e.g. Nadler et al., 2020; Fischer et al., 2024b). The DM self-interactions between host and satellite could be incorporated in simulations such as the ones we have presented following the approach by Zeng et al. (2022) or within the more general framework by Klemmer et al. (in prep.).
@@
Another physical effect that we neglect here is dynamical friction. However, given the low mass of our satellite halo compared to the host mass, it is expected to be hardly affected by dynamical friction. Nevertheless it would be possible to include the effect of dynamical friction without resolving the host system, but approximating the deceleration that the subhalo experiences (e.g. Petts et al., 2015).
@@
Our numerical scheme has the advantage that the pairwise interactions do not harm energy conservation, although we employ shared and distributed memory parallelisation. Many of the codes used today do not have this property. In our case it implies that the energy conservation is independent of the chosen kernel function for the SIDM computations, by construction. This is, because the value of the interaction probability or the drag force has no direct impact on the energy conservation. However, the chosen kernel function could eventually make a difference in the context of angular momentum conservation and for the effective heat conduction in the case of large values for $h/l$ (see Sect. 3.5).
@@
Ways to improve the accuracy of the SIDM simulations have already been discussed by Fischer et al. (2024b). We want to highlight that using a symmetric evaluation of the oct-tree used for the gravitational force computation would be helpful, e.g. as done by Appel (1985). Another aspect is the time-irreversible choice of the time step. Approaches that make the time stepping function more time symmetric can help to improve the accuracy, e.g. reduce the error on energy conservation (Dehnen, 2017).
@@
We also note that the above-mentioned strategies may not fully solve the challenges in simulating the late collapse phase. For example, it remains problematic to resolve smaller and smaller length scales while the halo is collapsing. Even if that would be resolved and also all numerical issues directly related to simulating SIDM, the modelling of gravity would still require a time step that continues shrinking while the halo is collapsing. The time step would become so small that the simulation can no longer be reasonably fast advanced in time. One way to overcome the challenges posed by the gravothermal collapse could be the introduction of a subgrid model that describes the very inner part of the halo. It might be a sink particle that accounts for the DM mass at the centre of the halo. In the post-collapse state, it may describe the black hole that eventually has formed and the DM spike that it is accreting. Idealised simulations of such a system can help to build such a model (Sabarish et al., 2025).
@@ 6 Conclusion
In this work, we have used the $N$-body simulation code OpenGaget3 to investigate the gravothermal collapse of SIDM halos. We studied the role of various numerical parameters and simulated the system in isolation and as a satellite halo. Additionally, we studied qualitative differences arising from the velocity and angular dependence of the self-interaction cross-section. Furthermore, the data for our highest-resolved simulation ($N=5\times 10^{7}$) is available [^2] The simulation data can be found at https://darkium.org. and may serve as a benchmark. Our main results and recommendations for SIDM simulations are as follows:
@@
    A very small softening length can lead to larger errors on the total energy because the time steps change more often in an asymmetric way. However, the softening length should be chosen small enough to resolve the relevant length scales, e.g. the size of the constant density core during collapse.
@@
    Overall, the halo evolution is very sensitive to energy conservation errors.
@@
    We did not find that a stricter gravity time step criterion for our SIDM simulation is needed, compared to what is commonly used for CDM simulations, i.e. $\eta=0.025$ is fine. At the same time, we may use a more stringent SIDM time step criterion than some other codes. It should be noted that simulations that evolve an SIDM halo over many more gravitational relaxation time scales than we did, may need a smaller value for $\eta$ (see also Sect. 4 by Fischer et al., 2024a).
@@
    Moreover, we recommend using a value for the opening criterion at least as small as $\alpha=5\times 10^{4}$, when a one-sided oct-tree is used.
@@
    A too large SIDM kernel size leads to an effectively too strong heat conduction, making the halo collapse faster. We find that $h/l$ should be below ten, but exceeding one by a few multiples appears to be unproblematic for the range that we have tested. This has implications on the number of particles required to resolve the collapse phase, i.e. it depends on how deep one wants to simulate in the gravothermal collapse.
@@
    We also want to point out that the maximum sampling radius of ICs has an influence on how fast the halo collapses. For a controlled simulation of an isolated NFW halo, we recommend using 15 $r_{s}$ instead of 10 $r_{s}$ for the truncation radius.
@@
    In our tests that enforce a minimum time step, we found that the central density increases further and stays constant at the same time the total energy artificially increases. While imposing a minimum time step in cosmological simulation may allow for correct predictions of the fraction of collapsed halos, we do not find evidence that the enclosed mass within a specific radius could be accurately predicted. Mass within larger radii ($\gtrsim r_{\mathrm{s}}$) might be overestimated.
@@
    The speed-up in halo evolution for a satellite compared to an isolated system is much larger for a velocity-dependent cross-section compared to a velocity-independent one, even when ignoring the scattering between the DM particles of the satellite with the host’s DM.
@@
Despite the numerical challenges, we have shown that it is possible to accurately simulate SIDM halos in the deep collapse phase in order to compare with observations. Taking the dense perturber of the GD-1 stellar stream as an example, our simulations can resolve the simulated halo within 10 pc, which is necessary to match the observations. It becomes prohibitively expensive to simulate extremely deep into the collapse phase using the $N$-body method, such as collapsing into a black hole. Nevertheless, in that regime, simulations based on schemes derived from first principle exploiting the symmetries of the problem can provide a complementary method (e.g. Gurian & May, 2025; Kamionkowski et al., 2025).
@@ Appendix A Overview of simulations
Table 1: Simulation properties and parameters.
Sim. $\sigma_{\mathrm{V}}/m$ $w$ angular $m_{\mathrm{n}}$ $N$ $r_{\mathrm{cut}}$ $\epsilon$ $\eta$ $N_{\mathrm{ngb}}$ $\tau$ $\Delta t_{\mathrm{min}}$
name $[\mathrm{cm}^{2}\mathrm{g}^{-1}]$ [km s-1] depend. [M] [$r_{\mathrm{s}}$] [pc] [Gyr]
@@
 [^3] The table gives the different simulation properties and parameters that we employ. The first column gives the simulation name with the colour as used in the figures. It follows the second column with the viscosity cross-section Eq. (10). The third column gives the velocity parameter of the velocity-dependent cross-section if applicable, in this case the second column corresponds to the normalisation of the cross-section, i.e. $\sigma_{0}$ of Eq. (4). The fourth column gives the angular dependence (iso. – isotropic, fwd. – forward dominated). The fifth column specifies whether the halo is evolved in an external potential, and the following columns give the numerical particle mass, the number of particles, the cut radius for sampling the ICs, the gravitational softening length ($\epsilon$), the gravitational time step parameter ($\eta$), the SIDM neighbour number $N_{\mathrm{ngb}}$, the SIDM time step parameter ($\tau$), and the lower limit of the allowed time step ($\Delta t_{\mathrm{min}}$). Moreover, the parameter for the tree node opening criterion is set to $\alpha=5\times 10^{-4}$ for all simulations following Fischer et al. (2024a). The simulations above the dashed line are evolved in isolation, and the ones below in an external potential motivated by the Milky Way as described in Sect. 4.1. The data for the highest-resolved SIDM simulation ($N=5\times 10^{7}$) is available at https://darkium.org.
With Tab. 1, we provide an overview of the simulations shown in this paper. Moreover, we use a consistent colour scheme for the figures with the colours as indicated in the table. The strength of the self-interactions is specified in terms of the viscosity cross-section. It is given by
@@ Appendix B The size of the time step
Figure 17: Variation of the gravitational time-step parameter $\eta$. In the upper panel, we show the mass enclosed within the inner 10 pc as a function of time. The lower panel gives the evolution of the total energy relative to the absolute initial value. The shown simulations are for the low resolution of $N=2\times 10^{6}$ and share the same parameters except $\eta$, which is varied as indicated in the legend. All parameters are given in Tab. 1 under the names H, I, and J. The shaded regions in the top panel indicate the uncertainty estimated based on shot noise.
Previous studies (Mace et al. 2024; Palubski et al. 2024) indicated that one may need to be careful in choosing the time step for simulations of SIDM halos to ensure accurate results. It has been claimed, that depending on the set-up, simulations of the SIDM halo collapse require more stringent settings for the time step criteria than in common CDM simulations. This has been particularly expressed in terms of choosing smaller values for the time step parameter of the gravitational time step criterion.
In our simulations, we find the value of the time step parameter $\eta$ has no significant influence on the results for the range that we have tested as shown in Fig. 17. We note that $\Delta t_{\mathrm{grav}}\propto\sqrt{\eta}$ (eq. 34 by Springel 2005). In conclusion, we do not need a stricter time step criterion for gravity compared to the CDM simulations. However, we also note that we use at the same time our time step criterion for SIDM (Fischer et al. 2024b). It becomes relevant at the late stage of the evolution ($t\gtrsim 4.4\,\mathrm{Gyr}$) for the particles in the inner region. Here, it results in smaller time steps than solely based on the gravitational time step criterion. Moreover, we want to point out that it might be important how exactly the SIDM time step criterion is formulated (see appendix B by Fischer et al. 2024b).
Furthermore, we want to point out that we have simulated a set-up for which the impact of the self-interactions is much stronger than numerical errors arising from solving for gravity. When simulating a DM halo over many more gravitational time scales it could be that the desired accuracy can only be achieved when choosing a smaller time step, i.e. a smaller value for $\eta$ (see also the discussion about the gravitational relaxation in Sect. 4 by Fischer et al. 2024a). However, if the numerical integration errors become significantly large, the collapse time of an SIDM halo simulation may no longer be inversely proportional to the cross-section. This might be in line with the results Mace et al. (2024) obtained for their simulations of small cross-sections that they have run for long time spans.

### v4

@@ 4.3 Angular and velocity dependence of the cross-section
Figure 15: Mass enclosed within 10 pc as a function of time for the satellite halos. The upper panel gives the results for a velocity-independent cross-section (simulations U, V, W, and X), and the lower panel is for velocity-dependent cross-sections (simulations Y and Z). More details on the simulations can be found in Table 1. In addition, the grey dotted line indicates the preferred mass range for the perturber of the stellar stream GD-1 (Bonaca et al., 2019).
In this section we investigate the role of the velocity and angular dependence of the self-interaction for the evolution of the satellite halo. As for the isolated halo, we employed isotropic and forward-dominated cross-sections that either follow the velocity dependence given by Eq. (4) or are velocity-independent. In detail, we employed the same cross-section as in Sect. 3.8. The results for the enclosed mass within 10 pc are shown in Fig. 15.
@@
Interestingly, we find that the angular dependence starts to play a role in the velocity-independent case, although the scatterings between the DM of the host and the satellite are neglected. The small-angle scattering leads here to a longer collapse time compared to the isotropic cross-section (Fig. 15). For the isolated halo, the two cross-sections resulted in the same collapse time (Fig. 12). Fischer et al. (2021a) show that the thermalisation process depends on the angular dependence of the cross-section, i.e. large-angle scattering populates the high-velocity tail of the Maxwell-Boltzmann distribution more quickly compared to small-angle scattering. In combination with the tidal forces, this allows us to explain the different collapse times. Moreover, the difference arising from the angular dependence is minor in the velocity-dependent case where scattering with a high relative velocity is strongly suppressed (bottom panel of Fig. 15). We also checked that both angular-dependent and -independent cases give rise to almost the same total gravitationally bound mass of the satellite halo regardless of the velocity dependence. For the runs shown here, the angular dependence of the scattering does not affect the global properties of the halo.
@@ 4.4 Fit with King model
Figure 16: Density profile fitted by a King model for a satellite halo. The results are for the simulation with a velocity-dependent isotropic cross-section. All simulation parameters are given in Table 1 under simulation Y.
@@
With this, we follow Zhang et al. (2025) who also used a King profile to describe the collapse phase of a potential GD-1 perturber, motivated by research on star clusters. To determine the density and radial parameters ($\rho_{0}$, $r_{\mathrm{c}}$), we used a limited radial range with $r<0.5\,\mathrm{kpc}$ only. For the fit, we maximised a likelihood based on Poisson statistics analogously to the description in Sect. 4 by Fischer et al. (2021a).
In Fig. 16, we show the fit of the King model for the collapse phase using our simulation with a velocity-dependent isotropic cross-section (simulation Y according to Table 1). It becomes visible that the King model provides a reasonable fit to the inner region of a satellite halo that undergoes core collapse. This is in line with the findings of Zhang et al. (2025) and may work well enough to fit the observational data, as discussed by the same authors. In addition, we fit the King mode to the isolated halo and show the density profile as well as the time evolution of the fitted parameter in Appendix D.
@@ 5 Discussion
In this section we discuss the limitations of our work and of the employed numerical scheme. We mention the physical processes that we neglected and that would be worth including in the modelling, and we highlight directions to improve the modelling of collapsing SIDM halos.
@@
In our simulations we neglected the scatterings between satellite and host particles. However, they can play a crucial role in shaping the evolution of the satellite. How strongly the DM of the satellite is affected depends on how strong the cross-section is on the relevant velocity scale, which is typically larger than for the velocities within the satellite. Hence, for a cross-section that is decreasing as a function of velocity, this is less important than for a velocity-independent cross-section (e.g. Silverman et al., 2022; Zeng et al., 2022, 2025b). In addition, the angular dependence of the cross-section can play an important role. Satellite halos are strongly affected if the cross-section is forward-dominated (Fischer et al., 2022) and the effect becomes weaker for velocity-dependent cross-sections (e.g. Nadler et al., 2020; Fischer et al., 2024b). The DM self-interactions between host and satellite could be incorporated in simulations such as the ones we have presented following the approach by Zeng et al. (2022) or within the more general framework of Klemmer et al. (in prep.).
@@
Another physical effect that we neglect here is dynamical friction. However, given the low mass of our satellite halo compared to the host mass, it is expected to be hardly affected by dynamical friction. Nevertheless, it would be possible to include the effect of dynamical friction without resolving the host system, but by approximating the deceleration experienced by the subhalo (e.g. Petts et al., 2015).
@@
Our numerical scheme has the advantage that the pairwise interactions do not harm energy conservation, although we employ shared and distributed memory parallelisation. Many of the codes used today do not have this property. In our case, it implies that the energy conservation is independent of the chosen kernel function for the SIDM computations, by construction. This is because the value of the interaction probability or the drag force does not have a direct impact on energy conservation. However, the chosen kernel function could eventually make a difference in the context of angular momentum conservation and for the effective heat conduction in the case of large values for $h/l$ (see Sect. 3.5).
@@
Ways to improve the accuracy of SIDM simulations have already been discussed by Fischer et al. (2024b). We want to highlight that using a symmetric evaluation of the oct-tree used for the gravitational force computation would be helpful (e.g. as done by Appel (1985)). Another aspect is the time-irreversible choice of the time step. Approaches that make the time-stepping function more time symmetric can help to improve the accuracy, for example by reducing the error on energy conservation (Dehnen, 2017).
@@
We also note that the above-mentioned strategies may not fully solve the challenges in simulating the late collapse phase. For example, it remains problematic to resolve smaller and smaller length scales while the halo is collapsing. Even if this were resolved, and all numerical issues directly related to simulating SIDM as well, the modelling of gravity would still require a time step that continues shrinking while the halo is collapsing. The time step would become so small that the simulation could no longer be reasonably fast advanced in time. One way to overcome the challenges posed by gravothermal collapse could be the introduction of a subgrid model that describes the very inner part of the halo. It might be a sink particle that accounts for the DM mass in the centre of the halo. In the post-collapse state, it may describe the black hole that finally formed and the DM spike that it is accreting. Idealised simulations of such a system can help build such a model (Sabarish et al., 2025).
@@ 6 Conclusion
In this work, we used the $N$-body simulation code OpenGadget3 to investigate the gravothermal collapse of SIDM halos. We studied the role of various numerical parameters and simulated the system in isolation and as a satellite halo. Additionally, we studied the qualitative differences arising from the velocity and angular dependence of the self-interaction cross-section. Furthermore, the data for our most highly resolved simulation ($N=5\times 10^{7}$) are available [^2] The simulation data can be found at https://darkium.org. and may serve as a benchmark. Our main results and recommendations for SIDM simulations are as follows:
@@
    A very small softening length can lead to larger errors in the total energy because the time steps change more often in an asymmetric way. However, the softening length chosen should be small enough to resolve the relevant length scales (e.g. the size of the constant density core during collapse).
@@
    Overall, halo evolution is very sensitive to energy conservation errors.
@@
    We did not find that a stricter gravity time step criterion is needed for our SIDM simulations compared to what is commonly used for CDM simulations, i.e. $\eta=0.025$ is fine. At the same time, we may use a more stringent SIDM time step criterion than some other codes. It should be noted that simulations that evolve an SIDM halo over many more gravitational relaxation timescales than we did may need a smaller value for $\eta$ (see also Sect. 4 by Fischer et al., 2024a).
@@
    Moreover, we recommend using a value for the opening criterion as small as $\alpha=5\times 10^{-4}$, when using a one-sided oct-tree.
@@
    A SIDM kernel size that is too large effectively leads to heat conduction that is too strong, which causes the halo to collapse faster. We find that $h/l$ should be below 10, but exceeding 1 by a few multiples appears to be unproblematic for the range that we have tested. This has implications on the number of particles required to resolve the collapse phase, i.e. it depends on how deep one wants to simulate in the gravothermal collapse.
@@
    We also note that the maximum sampling radius of ICs has an influence on how fast the halo collapses. For a controlled simulation of an isolated NFW halo, we recommend using 15 $r_{s}$ instead of 10 $r_{s}$ for the truncation radius.
@@
    In our tests that enforce a minimum time step, we found that the central density increases further and stays constant at the same time that the total energy artificially increases. Although imposing a minimum time step in cosmological simulations may allow correct predictions of the fraction of collapsed halos to be made, we do not find evidence that the enclosed mass within a specific radius can be accurately predicted. The mass within larger radii ($\gtrsim r_{\mathrm{s}}$) might be overestimated.
@@
    The speed-up in halo evolution for a satellite compared to an isolated system is much greater for a velocity-dependent cross-section compared to a velocity-independent one, even when ignoring the scattering between the DM particles of the satellite with the host’s DM.
@@
Despite the numerical challenges, we have shown that it is possible to accurately simulate SIDM halos in the deep-collapse phase to compare with observations. Taking the dense perturber of the GD-1 stellar stream as an example, our simulations can resolve the simulated halo within 10 pc, which is necessary to match the observations. It becomes prohibitively expensive to simulate extremely deep into the collapse phase using the $N$-body method, such as collapsing into a black hole. Nevertheless, in that regime, simulations based on schemes derived from first principles exploiting the symmetries of the problem can provide a complementary method (e.g. Gurian & May, 2025; Kamionkowski et al., 2025).
@@ Appendix A Overview of simulations
Table 1: Simulation properties and parameters. [^3] The table gives the different simulation properties and parameters that we employ. The first column gives the simulation name with the colour as used in the figures. It follows the second column with the viscosity cross-section Eq. (10). The third column gives the velocity parameter of the velocity-dependent cross-section if applicable, in this case the second column corresponds to the normalisation of the cross-section, i.e. $\sigma_{0}$ of Eq. (4). The fourth column gives the angular dependence (iso. – isotropic, fwd. – forward dominated). The fifth column specifies whether the halo is evolved in an external potential, and the following columns give the numerical particle mass, the number of particles, the cut radius for sampling the ICs, the gravitational softening length ($\epsilon$), the gravitational time step parameter ($\eta$), the SIDM neighbour number $N_{\mathrm{ngb}}$, the SIDM time step parameter ($\tau$), and the lower limit of the allowed time step ($\Delta t_{\mathrm{min}}$). Moreover, the parameter for the tree node opening criterion is set to $\alpha=5\times 10^{-4}$ for all simulations following Fischer et al. (2024a). The simulations above the dashed line are evolved in isolation, and the ones below in an external potential motivated by the Milky Way as described in Sect. 4.1. The data for the highest-resolved SIDM simulation ($N=5\times 10^{7}$) is available at https://darkium.org.
Sim. $\sigma_{\mathrm{V}}/m$ $w$ angular $m_{\mathrm{n}}$ NN $r_{\mathrm{cut}}$ $\epsilon$ $\eta$ $N_{\mathrm{ngb}}$ $\tau$ $\Delta t_{\mathrm{min}}$
name $\mathrm{cm}^{2}\mathrm{g}^{-1}] [$km s-1] depend. [M] $r_{\mathrm{s}}$] [pc] [Gyr]
@@
With Table 1, we provide an overview of the simulations shown in this paper. Moreover, we use a consistent colour scheme for the figures with the colours as indicated in the table. The strength of the self-interactions is specified in terms of the viscosity cross-section. It is given by
@@ Appendix B The size of the time step
Figure 17: Variation of the gravitational time step parameter $\eta$. In the upper panel, we show the mass enclosed within the inner 10 pc as a function of time. The lower panel gives the evolution of the total energy relative to the absolute initial value. The shown simulations are for the low resolution of $N=2\times 10^{6}$ and share the same parameters except $\eta$, which is varied as indicated in the legend. All parameters are given in Table 1 under the names H, I, and J. The shaded regions in the top panel indicate the uncertainty estimated based on shot noise.
Previous studies (Mace et al. 2024; Palubski et al. 2024) indicated that one may need to be careful in choosing the time step for simulations of SIDM halos to ensure accurate results. It has been claimed that depending on the set-up, simulations of the SIDM halo collapse require more stringent settings for the time step criteria than in common CDM simulations. This has been particularly expressed in terms of choosing smaller values for the time step parameter of the gravitational time step criterion.
In our simulations, we find that the value of the time step parameter $\eta$ has no significant influence on the results for the range that we have tested as shown in Fig. 17. We note that $\Delta t_{\mathrm{grav}}\propto\sqrt{\eta}$ (eq. 34 by Springel 2005). In conclusion, we do not need a stricter time step criterion for gravity compared to the CDM simulations. However, we also note that we use at the same time our time step criterion for SIDM (Fischer et al. 2024b). It becomes relevant at the late stage of the evolution ($t\gtrsim 4.4\,\mathrm{Gyr}$) for the particles in the inner region. Here, it results in smaller time steps than solely based on the gravitational time step criterion. Moreover, we want to point out that it might be important how exactly the SIDM time step criterion is formulated (see appendix B by Fischer et al. 2024b).
Furthermore, we want to point out that we have simulated a set-up for which the impact of the self-interactions is much stronger than numerical errors arising from solving for gravity. When simulating a DM halo over many more gravitational timescales, it could be that the desired accuracy can only be achieved when choosing a smaller time step, i.e. a smaller value for $\eta$ (see also the discussion about the gravitational relaxation in Sect. 4 by Fischer et al. 2024a). However, if the numerical integration errors become significantly large, the collapse time of an SIDM halo simulation may no longer be inversely proportional to the cross-section. This might be in line with the results Mace et al. (2024) obtained for their simulations of small cross-sections that they have run for long time spans.